recombination in phages

you have isolated two mutations that affect plaque morphology in phages (y and z). you conduct a genetic cross between the two phage strains by coinfecting bacteria with a wild-type strain and a strain that carries both mutations. phage chromosomes enter the host and replicate. you collect some of the progeny phages and plate them onto a lawn of bacterial cells. you observe the following numbers of plaques:

3042 480 513 513 2965 7000

1. which are examples of recombinant progeny?
Select all that apply.
y+z+, y+z, yz+, yz

2. what percentage of progeny were recombinant?_________%

3. what is the frequency of the recombination between the y and z genes? ___________ %

4. estimate the map distance between the y and z genes _________ map units

please help. I am totally lost.
Answer #1

Answer:

1). Hint: Always recombinant genotypes are less number than the non-recombinant genotypes. Hence, recombinant progeny are y+ z & y z+.

2).

Total recombinants = 480+513 = 993

Percentage of progeny were recombinant = (993/7000)100 = 14.19

3).

Recombination frequency (RF) between the y and z genes= (no. of recombinant progeny / total progeny)100

RF= (993/7000)100 = 14.19%

4).

RF(%) = Distance between genes (mu)

Distance between eye color and body color = 14.19 m.u.